# WaveCert: Formal Compiler Validation for Asynchronous Dataflow Programs

### v1

@@ 4.1. Describing a Cut-Bisimulation via Cut Points
To finitely describe a proposed cut-bisimulation, we use a list of pairs of dataflow and LLVM cut points. A cut point is a pair $P=(c_{\mathrm{df}}(\mathbf{x}),c_{\mathrm{llvm}}(\mathbf{y}))\mid\varphi(\mathbf{x},\mathbf{y})$ of symbolic dataflow and LLVM configurations constrained by the correspondence condition $\varphi(\mathbf{x},\mathbf{y})$. In our case, the condition $\varphi(\mathbf{x},\mathbf{y})$ is a conjunction of equations between variables in $\mathbf{x}$ and $\mathbf{y}$. In particular, we require $\mathrm{mem}(c_{\mathrm{df}}(\mathbf{x}))=\mathrm{mem}(c_{\mathrm{llvm}}(\mathbf{y}))$ to be in $\varphi(\mathbf{x},\mathbf{y})$ for any cut point. Semantically, each cut point describes the relation $\mathcal{R}(P):=\{(c_{\mathrm{df}}(\mathbf{x}),c_{\mathrm{llvm}}(\mathbf{y}))\in\mathcal{C}_{\mathrm{df}}\times\mathcal{C}_{\mathrm{llvm}}\mid\mathbf{x},\mathbf{y}\in V,\varphi(\mathbf{x},\mathbf{y})\}$. For a list of $n$ cut points
$\{P^{i}=(c^{i}_{\mathrm{df}}(\mathbf{x^{i}}),c^{i}_{\mathrm{llvm}}(\mathbf{y^{i}}))\mid\varphi^{i}(\mathbf{x^{i}},\mathbf{y^{i}})\}_{i=1}^{n},$
@@
To ensure that the cut-bisimulation we propose satisfies the property in Definition 2, we also require that the first cut point must be exactly the pairs of initial configurations with equal memory. That is $\mathcal{R}(P^{1})=\{(c_{\mathrm{df}},c_{\mathrm{llvm}})\in\mathcal{I}_{\mathrm{df}}\times\mathcal{I}_{\mathrm{llvm}}\mid\mathrm{mem}_{\mathrm{df}}(c_{\mathrm{df}})=\mathrm{mem}_{\mathrm{llvm}}(c_{\mathrm{llvm}})\}$. Furthermore, we require that the last cut point $P^{n}$ must be the pairs of final configurations with equal memory, that is $\mathcal{R}(P^{n})=\{(c_{\mathrm{df}},c_{\mathrm{llvm}})\in\mathcal{C}_{\mathrm{df}}\times\mathcal{C}_{\mathrm{llvm}}\mid c_{\mathrm{df}},c_{\mathrm{llvm}}\text{ final},\mathrm{mem}_{\mathrm{df}}(c_{\mathrm{df}})=\mathrm{mem}_{\mathrm{llvm}}(c_{\mathrm{llvm}})\}$. Then $\mathcal{R}(P^{1},\dots,P^{n})$ satisfies Definition 2.
4.2. Checking Cut-Bisimulation
Now that we have a way to symbolically describe a relation $\mathcal{R}(P^{1},\dots,P^{n})$ with cut points satisfying Definition 2, we need to check that it is indeed a cut-bisimulation.
To achieve this, we perform symbolic execution from each cut point and try to reach another cut point. Suppose we are checking cut point $P^{i}=(c^{i}_{\mathrm{df}}(\mathbf{x^{i}}),c^{i}_{\mathrm{llvm}}(\mathbf{y^{i}}))\mid\varphi^{i}(\mathbf{x^{i}},\mathbf{y^{i}})$. We perform symbolic execution from $c^{i}_{\mathrm{df}}(\mathbf{x^{i}})$ and $c^{i}_{\mathrm{llvm}}(\mathbf{y^{i}})$ until all of their symbolic branches hit another cut point. For the LLVM side, we get a list of $k_{i}$ symbolic branches with path conditions denoted by $\psi$:
$c^{i}_{\mathrm{llvm}}(\mathbf{y^{i}})\to^{+}_{\mathrm{llvm}}c^{i,1}_{\mathrm{llvm}}(\mathbf{y^{i}})\mid\psi_{\mathrm{llvm},1}(\mathbf{y^{i}})$
@@ 4.2. Checking Cut-Bisimulation
$c^{i}_{\mathrm{llvm}}(\mathbf{y^{i}})\to^{+}_{\mathrm{llvm}}c^{i,k_{i}}_{\mathrm{llvm}}(\mathbf{y^{i}})\mid\psi_{\mathrm{llvm},k_{i}}(\mathbf{y^{i}})$
where each right hand side $c^{i,l}_{\mathrm{llvm}}(\mathbf{y^{i}})$ for $l\in\{1,\dots,k_{i}\}$ is an instance $c^{j}_{\mathrm{llvm}}(\sigma)$ of some other LLVM cut point $c^{j}_{\mathrm{llvm}}(\mathbf{y^{j}})$ with substitution $\sigma:\mathbf{y^{j}}\to V(\mathbf{y^{i}})$. Furthermore, since our LLVM semantics is deterministic, the path conditions should be disjoint, i.e., for any $l_{1},l_{2}\in\{1,\dots,k_{i}\}$, if $l_{1}\neq l_{2}$, $\psi_{\mathrm{llvm},l_{1}}(\mathbf{y^{i}})\wedge\psi_{\mathrm{llvm},l_{2}}(\mathbf{y^{i}})$ is unsatisfiable.
For the dataflow cut point $c^{i}_{\mathrm{df}}(\mathbf{x^{i}})$, since the dataflow semantics is nondeterministic, we try to replicate the execution of LLVM branches on the dataflow side. For each LLVM branch $l\in\{1,\dots,k_{i}\}$, we can get a trace of LLVM instructions that are executed from $c^{i}_{\mathrm{llvm}}(\mathbf{y^{i}})$ to $c^{i,l}_{\mathrm{llvm}}(\mathbf{y^{i}})$. Using hints from the compiler, we know that a subset of these LLVM instructions map to a list of dataflow operators $(o^{l}_{1},\dots,o^{l}_{n_{l}})$. For each LLVM branch, to produce the corresponding dataflow branch, we symbolically execute the dataflow cut point $c^{i}_{\mathrm{df}}(\mathbf{x^{i}})$ on the schedule $(o^{l}_{1},\dots,o^{l}_{n_{l}})$ with the path condition $\exists\mathbf{y^{i}}\varphi^{i}(\mathbf{x^{i}},\mathbf{y^{i}})\wedge\psi_{\mathrm{llvm},l}(\mathbf{y^{i}})$:
$c^{i}_{\mathrm{df}}(\mathbf{x^{i}})\to^{o^{l}_{1}*}_{\mathrm{df}}\dots\to^{o^{l}_{n_{l}}*}_{\mathrm{df}}c^{i,l}_{\mathrm{df}}(\mathbf{x^{i}})\mid\exists\mathbf{y^{i}}\varphi^{i}(\mathbf{x^{i}},\mathbf{y^{i}})\wedge\psi_{\mathrm{llvm},l}(\mathbf{y^{i}})$
where each $\to^{o*}_{\mathrm{df}}$ denotes the dataflow transition by first firing the operator $o$, and then firing any available steer operator until none is fireable anymore. The canonical schedule $\to_{\mathrm{df\mid canon}}$ is defined by this order of execution of operators.
Therefore, by replicating the execution of the LLVM branches from cut point $i$, we get a list of dataflow branches:
(1) $c^{i}_{\mathrm{df}}(\mathbf{x^{i}})\to^{+}_{\mathrm{df\mid canon}}c^{i,1}_{\mathrm{df}}(\mathbf{x^{i}})\mid\psi_{\mathrm{df},1}(\mathbf{x^{i}})$
@@
$c^{i}_{\mathrm{df}}(\mathbf{x^{i}})\to^{+}_{\mathrm{df\mid canon}}c^{i,k_{i}}_{\mathrm{df}}(\mathbf{x^{i}})\mid\psi_{\mathrm{df},k_{i}}(\mathbf{x^{i}})$
where each path condition $\psi_{\mathrm{df},l}(\mathbf{x^{i}}):=\exists\mathbf{y^{i}}\varphi^{i}(\mathbf{x^{i}},\mathbf{y^{i}})\wedge\psi_{\mathrm{llvm},l}(\mathbf{y^{i}})$. We check that the dataflow symbolic execution from cut point $i$ has covered all concrete cases, i.e., all possible cases of $\mathbf{x^{i}}$, by checking the unsatisfiability of $\neg\bigvee^{k_{i}}_{l=1}\psi_{\mathrm{df},l}(\mathbf{x^{i}})$.
Finally, we check that each pair of dataflow and LLVM branches $(c^{i,l}_{\mathrm{df}}(\mathbf{x^{i}}),c^{i,l}_{\mathrm{llvm}}(\mathbf{y^{i}}))$ matches some target cut point $P^{j}=(c^{j}_{\mathrm{df}}(\mathbf{x^{j}}),c^{j}_{\mathrm{llvm}}(\mathbf{y^{j}}))\mid\varphi^{j}(\mathbf{x^{j}},\mathbf{y^{j}})$:
$\{(c^{i,l}_{\mathrm{df}}(\mathbf{x^{i}}),c^{i,l}_{\mathrm{llvm}}(\mathbf{y^{i}}))\mid\varphi^{i}(\mathbf{x^{i}},\mathbf{y^{i}})\wedge\psi_{\mathrm{df},l}(\mathbf{x^{i}})\wedge\psi_{\mathrm{llvm},l}(\mathbf{y^{i}})\}\subseteq\mathcal{R}(P^{j})$
or equivalently, there are substitutions $\sigma^{i,l}_{\mathrm{df}}:\mathbf{x^{j}}\to V(\mathbf{x^{i}}),\sigma^{i,l}_{\mathrm{llvm}}:\mathbf{y^{j}}\to V(\mathbf{y^{i}})$ such that the following conditions hold:
@@
  • The LLVM branch is an instance of the target LLVM cut point: $c^{i,l}_{\mathrm{llvm}}(\mathbf{y^{i}})=c^{j}_{\mathrm{llvm}}(\sigma^{i,l}_{\mathrm{llvm}})$.
@@
    $\varphi^{i}(\mathbf{x^{i}},\mathbf{y^{i}})\wedge\psi_{\mathrm{df},l}(\mathbf{x^{i}})\wedge\psi_{\mathrm{llvm},l}(\mathbf{y^{i}})\rightarrow\varphi^{j}(\sigma^{i,l}_{\mathrm{df}},\sigma^{i,l}_{\mathrm{llvm}})$
If the check described in this section is passed for all cut points $i$ and all branches $1,\dots,k_{i}$, we can conclude that $\mathcal{R}(P^{1},\dots,P^{n})$ is indeed a cut-bisimulation.
@@ 4.3. Heuristics to Infer Cut Points
The final step is to actually infer the LLVM and dataflow cut points. In general, program equivalence for Turing-complete models like LLVM and dataflow programs is undecidable (rogers-87, ), so our best hope is to have good heuristics to infer the cut points for the specific input/output programs of the compiler.
The placement of cut points is important for the symbolic execution to terminate. It is in theory correct (for terminating dataflow and LLVM programs) to have exactly two cut points: one at the initial configurations, and one for all final configurations. But the symbolic execution will try to simulate all possible traces together, which are unbounded, and the check may not terminate. To ensure termination, on the LLVM side, we place a cut point at each of the following program points:
  • The entry of the function (i.e. the initial configuration);
  • The back edge of each loop header block; and
  • The exit program point.
For the dataflow program, we need to place the first cut point at the initial configuration of the dataflow program, and the last cut point at the final configuration. For loops, ideally, we want to place the cut points at the corresponding “program points” at loop headers. However, due to the lack of control-flow in the dataflow program, the form of the corresponding dataflow configurations at these “program points” is hard to infer statically. Instead of generating the dataflow cut points ahead of time, we dynamically infer them during symbolic execution of dataflow branches. For example, after mirroring the LLVM branch execution in a dataflow branch such as in Equation 1, if the dataflow cut point corresponding to the LLVM cut point has not been inferred yet, we generalize the symbolic configuration on the RHS of the dataflow branch Equation 1 to a dataflow cut point by replacing symbolic expressions with fresh variables. The correspondence condition is then inferred using compiler hints to match fresh variables in the new dataflow cut point to an LLVM variable defined by some LLVM instruction.
Note that we do not need these heuristics to be sound. As long as the cut points for the initial and final dataflow/LLVM configurations are generated correctly and the cut-bisimulation check in Section 4.2 succeeds, the soundness of the cut-bisimulation follows. While we do not prove these heuristics correct, they work well in practice and cover all test cases in our evaluation (Section 6).

### v2

@@ 4.2. Describing a Cut-Simulation via Cut Points
To finitely describe a proposed cut-simulation, we use a list of pairs of dataflow and Seq cut points. A cut point is a pair $P=(c_{\mathrm{df}}(\mathbf{x}),c_{\mathrm{seq}}(\mathbf{y}))\mid\varphi(\mathbf{x},\mathbf{y})$ of symbolic dataflow and Seq configurations constrained by the correspondence condition $\varphi(\mathbf{x},\mathbf{y})$. In our case, the condition $\varphi(\mathbf{x},\mathbf{y})$ is a conjunction of equalities between variables in $\mathbf{x}$ and $\mathbf{y}$.
Semantically, each cut point describes the binary relation $\mathcal{R}(P):=\{(c_{\mathrm{df}}(\mathbf{x}),c_{\mathrm{seq}}(\mathbf{y}))\in\mathcal{C}_{\mathrm{df}}\times\mathcal{C}_{\mathrm{seq}}\mid\mathbf{x},\mathbf{y}\in V,\varphi(\mathbf{x},\mathbf{y})\}$. For a list of $n$ cut points
$\{P^{i}=(c^{i}_{\mathrm{df}}(\mathbf{x^{i}}),c^{i}_{\mathrm{seq}}(\mathbf{y^{i}}))\mid\varphi^{i}(\mathbf{x^{i}},\mathbf{y^{i}})\}_{i=1}^{n},$
@@
To ensure that the cut-simulation we propose satisfies the three conditions in Definition 6, we only construct cut points such that:
  • For any cut point $P^{i}$, the correspondence constraint $\varphi^{i}(\mathbf{x}^{i},\mathbf{y}^{i})$ implies equal memory $\mathsf{Mem}(c^{i}_{\mathrm{df}}(\mathbf{x}^{i}))=\mathsf{Mem}(c^{i}_{\mathrm{seq}}(\mathbf{y}^{i}))$.
  • The first cut point exactly represents the pairs of equivalent, initial configurations with equal memory; i.e., $\mathcal{R}(P^{1})=\{(c_{\mathrm{df}},c_{\mathrm{seq}})\in\ \sim_{\mathcal{I}}\ \mid\mathsf{Mem}(c_{\mathrm{df}})=\mathsf{Mem}(c_{\mathrm{seq}})\}$.
  • The last cut point $P^{n}$ exactly represents the pairs of final configurations with equal memory; i.e., $\mathcal{R}(P^{n})=\{(c_{\mathrm{df}},c_{\mathrm{seq}})\in\mathcal{C}_{\mathrm{df}}\times\mathcal{C}_{\mathrm{seq}}\mid c_{\mathrm{df}},c_{\mathrm{seq}}\text{ final},\mathsf{Mem}(c_{\mathrm{df}})=\mathsf{Mem}(c_{\mathrm{seq}})\}$.
Then $\mathcal{R}(P^{1},\dots,P^{n})$ satisfies Definition 6.
4.3. Checking Cut-Simulation
Now that we have a way to symbolically describe a relation $\mathcal{R}(P^{1},\dots,P^{n})$ with cut points satisfying Definition 6, we need to check that it is indeed a cut-simulation.
To achieve this, we perform symbolic execution from each cut point and try to reach another cut point. Suppose we are checking cut point $P^{i}=(c^{i}_{\mathrm{df}}(\mathbf{x^{i}}),c^{i}_{\mathrm{seq}}(\mathbf{y^{i}}))\mid\varphi^{i}(\mathbf{x^{i}},\mathbf{y^{i}})$. We perform the following three steps.
Step 1. Symbolic execution of Seq cut points
We perform symbolic execution from $c^{i}_{\mathrm{seq}}(\mathbf{y^{i}})$ and get a list of $k_{i}$ symbolic branches with path conditions denoted by $\psi$:
$c^{i}_{\mathrm{seq}}(\mathbf{y^{i}})\to^{+}_{\mathrm{seq}}c^{i,1}_{\mathrm{seq}}(\mathbf{y^{i}})\mid\psi_{\mathrm{seq},1}(\mathbf{y^{i}})$
@@ 4.3. Checking Cut-Simulation
$c^{i}_{\mathrm{seq}}(\mathbf{y^{i}})\to^{+}_{\mathrm{seq}}c^{i,k_{i}}_{\mathrm{seq}}(\mathbf{y^{i}})\mid\psi_{\mathrm{seq},k_{i}}(\mathbf{y^{i}})$
where each right hand side $c^{i,l}_{\mathrm{seq}}(\mathbf{y^{i}})$ for $l\in\{1,\dots,k_{i}\}$ is an instance $c^{j}_{\mathrm{seq}}(\sigma)$ of some Seq cut point $c^{j}_{\mathrm{seq}}(\mathbf{y^{j}})$ with substitution $\sigma:\mathbf{y^{j}}\to\mathsf{Terms}(\mathbf{y^{i}})$. That is, at the end of each branch, the configuration is matched to some cut point.
Note that since the initial symbolic configuration $c^{i}_{\mathrm{seq}}(\mathbf{y^{i}})$ does not have any constraints, the final path conditions should cover all cases of $\mathbf{y^{i}}$; in other words, $\bigvee^{k_{i}}_{l=1}\psi_{\mathrm{seq},l}(\mathbf{y^{i}})$ is valid.
Step 2. Symbolic execution of dataflow cut points on the canonical schedule
For the dataflow cut point $c^{i}_{\mathrm{df}}(\mathbf{x^{i}})$, since the dataflow semantics is nondeterministic, the choice of operators to fire is unclear. Therefore, we follow a canonical schedule $\mathsf{Canon}$ (Definition 3) to replicate the execution of Seq branches on the dataflow side.
For each Seq branch $l\in\{1,\dots,k_{i}\}$, we have a Seq trace $\tau_{\mathrm{seq}}=c^{i}_{\mathrm{seq}}(\mathbf{y}^{i})\dots c^{i,l}_{\mathrm{seq}}(\mathbf{y}^{i})$. By mapping this trace through the canonical schedule $\tau_{\mathrm{df}}:=\mathsf{Canon}(c^{i}_{\mathrm{df}}(\mathbf{x^{i}}),\tau_{\mathrm{seq}})$, we get a list of dataflow operators $(o^{l}_{1},\dots,o^{l}_{n_{l}})$ fired in $\tau_{\mathrm{df}}$. Then we symbolically execute the dataflow cut point $c^{i}_{\mathrm{df}}(\mathbf{x^{i}})$ on the schedule $(o^{l}_{1},\dots,o^{l}_{n_{l}})$ with the path condition $\psi_{\mathrm{df},l}(\mathbf{x^{i}}):=\exists\mathbf{y^{i}}\varphi^{i}(\mathbf{x^{i}},\mathbf{y^{i}})\wedge\psi_{\mathrm{seq},l}(\mathbf{y^{i}})$:
$c^{i}_{\mathrm{df}}(\mathbf{x^{i}})\to^{o^{l}_{1}}_{\mathrm{df}}\dots\to^{o^{l}_{n_{l}}}_{\mathrm{df}}c^{i,l}_{\mathrm{df}}(\mathbf{x^{i}})\mid\psi_{\mathrm{df},l}(\mathbf{x^{i}})$
The new path condition $\psi_{\mathrm{df},l}(\mathbf{x^{i}})$ amounts to replacing variables in the Seq path condition $\psi_{\mathrm{seq},l}(\mathbf{y^{i}})$ with corresponding variables in the correspondence constraint $\varphi^{i}(\mathbf{x^{i}},\mathbf{y^{i}})$. With the additional path condition imposed, the symbolic execution is not likely to branch. If it does, the cut-simulation check will fail; in which case, a more precise heuristic is needed for the canonical schedule.
By doing this for each Seq branch, we get a list of dataflow branches:
(1) $c^{i}_{\mathrm{df}}(\mathbf{x^{i}})\to^{+}_{\mathrm{df}}c^{i,1}_{\mathrm{df}}(\mathbf{x^{i}})\mid\psi_{\mathrm{df},1}(\mathbf{x^{i}})$
@@
$c^{i}_{\mathrm{df}}(\mathbf{x^{i}})\to^{+}_{\mathrm{df}}c^{i,k_{i}}_{\mathrm{df}}(\mathbf{x^{i}})\mid\psi_{\mathrm{df},k_{i}}(\mathbf{x^{i}})$
In our case, these branches should cover all possible values of $\mathbf{x}^{i}$ (i.e., $\bigvee^{k_{i}}_{l=1}\psi_{\mathrm{df},l}(\mathbf{x^{i}})$ is valid), since in the path condition $\psi_{\mathrm{df},l}(\mathbf{x^{i}})$, the correspondence constraint $\varphi^{i}(\mathbf{x^{i}},\mathbf{y^{i}})$ only contains a conjunction of equalities between variables in $\mathbf{x^{i}}$ and $\mathbf{y^{i}}$, and the union of all Seq path conditions $\bigvee^{k_{i}}_{l=1}\psi_{\mathrm{seq},l}(\mathbf{y^{i}})$ is valid. In general, however, if the correspondence constraint $\varphi^{i}(\mathbf{x^{i}},\mathbf{y^{i}})$ is more complex (e.g. containing formulas such as $x_{1}=2y_{2}$), we may need to additionally verify that $\bigvee^{k_{i}}_{l=1}\psi_{\mathrm{df},l}(\mathbf{x^{i}})$ is valid to ensure that all possible concrete configurations represented by $c^{i}_{\mathrm{df}}(\mathbf{x^{i}})$ is covered in the symbolic execution.
Step 3. Cut point subsumption
Finally, we check that each pair of dataflow and Seq branches $(c^{i,l}_{\mathrm{df}}(\mathbf{x^{i}}),c^{i,l}_{\mathrm{seq}}(\mathbf{y^{i}}))$ is contained in some target cut point $P^{j}=(c^{j}_{\mathrm{df}}(\mathbf{x^{j}}),c^{j}_{\mathrm{seq}}(\mathbf{y^{j}}))\mid\varphi^{j}(\mathbf{x^{j}},\mathbf{y^{j}})$:
(2) $\{(c^{i,l}_{\mathrm{df}}(\mathbf{x^{i}}),c^{i,l}_{\mathrm{seq}}(\mathbf{y^{i}}))\mid\varphi^{i}(\mathbf{x^{i}},\mathbf{y^{i}})\wedge\psi_{\mathrm{df},l}(\mathbf{x^{i}})\wedge\psi_{\mathrm{seq},l}(\mathbf{y^{i}})\}\subseteq\mathcal{R}(P^{j})$
or equivalently, there are substitutions $\sigma^{i,l}_{\mathrm{df}}:\mathbf{x^{j}}\to\mathsf{Terms}(\mathbf{x^{i}}),\sigma^{i,l}_{\mathrm{seq}}:\mathbf{y^{j}}\to\mathsf{Terms}(\mathbf{y^{i}})$ such that the following conditions hold:
@@
  • The Seq branch is an instance of the target Seq cut point: $c^{i,l}_{\mathrm{seq}}(\mathbf{y^{i}})=c^{j}_{\mathrm{seq}}(\sigma^{i,l}_{\mathrm{seq}})$.
@@
    $\varphi^{i}(\mathbf{x^{i}},\mathbf{y^{i}})\wedge\psi_{\mathrm{df},l}(\mathbf{x^{i}})\wedge\psi_{\mathrm{seq},l}(\mathbf{y^{i}})\rightarrow\varphi^{j}(\sigma^{i,l}_{\mathrm{df}},\sigma^{i,l}_{\mathrm{seq}})$
If the steps in this section are passed for all cut points $i$, we can conclude that $\mathcal{R}(P^{1},\dots,P^{n})$ is indeed a cut-simulation.
@@ 4.4. Heuristics to Infer Cut Points
In this section, we describe our heuristics to infer cut points specific to LLVM and dataflow programs. In general, program equivalence for Turing-complete models like LLVM and dataflow programs is undecidable, so our best hope is to have good heuristics to infer the cut points for the specific input/output programs of the compiler.
The placement of cut points is important for the symbolic execution to terminate. It is in theory correct (for terminating dataflow and LLVM programs) to have exactly two cut points: one at the initial configurations, and one for all final configurations. But the symbolic execution will try to simulate all possible traces together, which are unbounded, and the check may not terminate. To ensure termination, on the LLVM side, we place a cut point at the entry of the function (i.e., the initial configuration), the back edge of each loop header block, and the exit program point.
For the dataflow program, we need to place the first cut point at the initial configuration of the dataflow program, and the last cut point at the final configuration. For loops, ideally, we want to place the cut points at the corresponding “program points” at loop headers. However, due to the lack of control-flow in the dataflow program, the form of the corresponding dataflow configurations at these “program points” is hard to infer statically. Instead of generating the dataflow cut points ahead of time, we dynamically infer them during the symbolic execution of dataflow branches. For example, after mirroring the LLVM branch execution in a dataflow branch such as in Equation 1, if the dataflow cut point corresponding to the LLVM cut point has not been inferred yet, we generalize the symbolic configuration on the RHS of the dataflow branch Equation 1 to a dataflow cut point by replacing symbolic expressions with fresh variables. The correspondence condition is then inferred using compiler hints to match fresh variables in the new dataflow cut point to an LLVM variable defined by some LLVM instruction.
Note that we do not need these heuristics to be sound. As long as the cut points for the initial and final dataflow/LLVM configurations are generated correctly and the cut-simulation check in Section 4.3 succeeds, the soundness of the cut-simulation follows. While we do not prove that these heuristics are complete, they work well in practice and cover all test cases in our evaluation (Section 6).
If the heuristics generate incorrect cut points, the simulation check would still terminate but fail. To see this, on the LLVM side, our placement of the cut points guarantees termination because we place one at each loop header. On the dataflow side, since the execution is mirroring the LLVM execution and no additional search is done, the algorithm would terminate but may fail to match the resulting dataflow configuration against the cut point.